Evaluate ∫c 𝐹. π‘‘π‘Ÿ,where 𝐹 = 𝑋^2 βˆ’ π‘Œ^2𝑖 + π‘₯𝑦𝑗 and curve 𝐢 is the arc of the curve 𝑦 = 𝑋^3 from (0,0) to (2,8).


1
Expert's answer
2022-03-21T02:04:29-0400

"\\int_C{Fdr}=\\left[ \\begin{array}{c}\ty=x^3\\\\\tdy=3x^2dx\\\\\tx:0\\rightarrow 2\\\\\tF_1=x^2-y^2=x^2-x^6\\\\\tF_2=xy=x^4\\\\\\end{array} \\right] =\\int_0^2{\\left( x^2-x^6 \\right) dx+x^4\\cdot 3x^2dx}=\\\\=\\left( \\frac{x^3}{3}+\\frac{2x^7}{7} \\right) |_{0}^{2}=\\frac{824}{21}"